OpenStudy (anonymous):

evaluate the integral of 2x^2-4x/x^3-2x^2 dx

OpenStudy (psymon):

Actually might be simpler than you realize. Simply factor the numerator and denominator. 2x^2 - 4x factors into 2x(x-2) while the denominator factors to x^2(x-2). Therefore you'll see the (x-2) factors will cancel leaving you with integral 2x/x^2 dx. From there it should be pretty simple?

OpenStudy (anonymous):

did you try using partial fractions?

OpenStudy (anonymous):

nevermind, i totally missed the comment by Psymon. He explained it pretty well.

OpenStudy (anonymous):

its ok but thank you anyways

OpenStudy (anonymous):

did you figure it out?

OpenStudy (anonymous):

\[\frac{2 x^2-4 x}{x^3-2 x^2}=\frac{2}{x} \]